OpenStudy (anonymous):

A 925 kg car moving with a velocity of + 20.0 m/s collides with a stationary truck of unknown mass. The vehicles lo ck together and move off with a velocity of +6.75 m/s. What was the mass of the truck?

OpenStudy (michele_laino):

here we have to apply the total momentum conservation law. So we can write this: \[\Large {M_1}V = \left( {{M_1} + {M_2}} \right)v\]

OpenStudy (michele_laino):

where M_1 is the mass of the moving truck, M_2 is the mass of the truck at the rest, V_1 is the velocity of the first truck, and v is the final velocity of both trucks

OpenStudy (michele_laino):

oop's.. V is the initial velocity or the velocity of the moving truck

OpenStudy (michele_laino):

then we can write: \[\Large {M_2} = \frac{{{M_1}V}}{v} - {M_1} = \frac{{925 \times 20}}{{6.75}} - 925 = ...{\text{Kg}}\]